OpenStudy (anonymous):

I am not understing how my teacher came up with the answer 1/36 out of 1/9-1/4. I get 5/36

OpenStudy (amistre64):

how did you arrive at your conclusion?

OpenStudy (amistre64):

\[\frac{1}{9}-\frac14=n\] \[\frac{1(9)}{9}-\frac{1(9)}{4}=(9)n\] \[1-\frac{9}{4}=9n\] \[1(4)-\frac{9(4)}{4}=9(4)n\] \[4-9=36n\] \[-5=36n\] \[\frac{-5}{36}=n\]

OpenStudy (anonymous):

(1/9)*(1/4)=1/36

OpenStudy (anonymous):

I changed the LCD on both to 26 times them by the lcd and minused them? 1/9 =4/36 and 1/4=9/36

OpenStudy (amistre64):

i never bother trying to determine the commons; i let the math do it for me

OpenStudy (anonymous):

Ok, I got it turns in to a multilpy problem - Thank you everyone!

OpenStudy (amistre64):

did you mistype the original then?

OpenStudy (amistre64):

1/9-1/4 does not "turn into a multiply problem". If you typed up the wrong problem, then yes; any number of outcomes might satisify ....

OpenStudy (stormfire1):

@amistre64: How did you the fractions bars in there? I've been trying to figure that out...

OpenStudy (amistre64):

the latex coding is: \frac{top}{bottom}

OpenStudy (amistre64):

but you have to wrap it in delimiters

OpenStudy (stormfire1):

I see..cool, thx :)

OpenStudy (amistre64):